Calculate the volume occupied by 22.12 g of sodium phosphate at 6.54 atm and 76 degrees Celsius.


1
Expert's answer
2022-04-06T14:24:02-0400

Since the substance is not gaseous at the conditions provided, we use density to find volume;

Density "=\dfrac{mass}{volume}"


Thus volume = "\dfrac{mass}{density}" "=\dfrac{22.12g}{1.62g/cm^3}" = 13.65cm3